OpenStudy (anonymous):

^3square root ot 27(x+6)^3

OpenStudy (anonymous):



OpenStudy (anonymous):

the 3rd root of 27 is just 3 and 3 root of (x+6)^3 = (x+6) so you get 3(x+6) = 3x + 18

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

np glad to help